# PASTA: A Scalable Framework for Multi-Policy AI Compliance Evaluation

### v1

@@ 4.6. Cost Optimization Strategies: Policy Chunking and Irrelevancy Filtering
Table 4. Empirical results showing how irrelevancy filtering reduces the number of comparison pairs and total runtime. The full evaluation compares all 23 model card sections against five policies—AIDA (40 articles), CCPA (54 articles), EU AI Act (113 articles), GDPR (99 articles), and the Colorado AI Act (8 articles)—resulting in 4,907 initial section–article pairs. Irrelevancy filtering removes pairs that consistently provide no contribution to compliance assessment, reducing the evaluation volume by 44.9% without loss of coverage.
@@
To support Design Goal 3, PASTA aggregates detailed pairwise evaluations into summaries that make large-scale compliance results easy to interpret. The evaluation process described in Section  4.5 generates a structured dataset consisting of one evaluation result for each comparison pair. Each result includes the corresponding model card section, policy article number, violation score, and an accompanying rationale. An example of the output is provided in Appendix. 10. The results are formatted as Markdown tables and are subsequently aggregated by the LLM into interpretable compliance summaries and actionable guidance. By condensing the detailed results into targeted summaries, the system helps users efficiently understand overall compliance status, identify policy-specific gaps, and highlight system components requiring attention. The evaluation results are aggregated into the following summaries:
  • Overall Summary: Provides a holistic assessment of the AI system’s compliance across all selected policies (see interface in Fig. 8 in Appendix.  A.6).
  • Policy-wise Summary: Presents a focused evaluation of the system’s compliance with individual policies, offering policy-specific insights (see interface in Fig. 9 in Appendix.  A.6).
@@
Figure 5. Section-wise Evaluation Interface linking compliance issues to specific model card sections. Each row presents the user’s documentation, detected issues, and concrete recommendations for remediation, supporting actionable system improvements. This table also supports policy-specific filtering
@@ 5.1.1. Expert Recruitment
We recruited 3 AI–policy compliance experts from Upwork, a globally used freelancing platform commonly relied upon in academic research to source verified professional contractors with specialized expertise. Experts were selected based on demonstrated experience interpreting “AI regulatory frameworks” and they were paid \$50 USD per hour. The 3 recruited experts reflected wide diversity in legal training. They were licensed attorneys, holding JD, LLB, and LLM degrees alongside jurisdiction-specific bar qualifications, with 6–10 years of legal practice across jurisdictions including the EU, North America, Latin America, the UK, and the Middle East. Their professional backgrounds covered privacy and data protection, AI governance and regulation, technology and cyberlaw, and commercial or corporate compliance. Each expert had prior experience assessing AI-related documentation and was familiar with major regimes such as the EU AI Act and CCPA.
@@ 5.1.2. Sampling of Policies and Model Cards
We sampled 18 policy articles across the EU AI Act and CCPA to ensure thematic breadth and balanced coverage, and sampled 3 diverse model cards among the 8 used in the preprocessing procedure detailed in Section 4.4 . To identify high-salience model card sections, we ran PASTA across all model-card sections and selected the 6 with the highest average violation scores across both policies. This resulted in 18 model-card sections which, when paired with the 18 curated articles, yielded 324 section–policy comparison pairs as our evaluation set. A full list of selected policy articles is provided in Appendix  B.
@@ 5.2. Findings
Overall, the results indicate that PASTA offer judgments that are comparably stable and human-aligned.
Before evaluating PASTA, we first assessed whether the 3 human experts formed a reliable evaluation baseline. Their intraclass correlation coefficients were .5940 for evaluating violation score and .6393 for relevance score, which can be interpreted as moderate-to-strong agreement based on established ICC guidelines  (Braun and Clarke, 2006; Koo and Li, 2016). This confirms that the expert-annotated dataset provides a consistent foundation for model comparison.
@@
PASTA’s compliance reasoning closely align with those of the human experts. Its violation and relevance scores closely mirrored expert results, reflected in Spearman correlations of $\rho=.626$ for violation scores and $\rho=.7611$ for relevance scores (see Table 5). Using widely cited interpretation thresholds (Cohen: .50–.69 = large; Mukaka: .60–.79 = strong), these values represent strong correlations  (Cohen, 1988; Mukaka, 2012). Its score magnitudes remained well-calibrated: mean absolute error was low (.284.284 for violation; .542.542 for relevance score), and the system matched expert scores within one point in 94.14% and 87.35% of cases respectively. These patterns indicate that PASTA approximate score levels in ways that align with expert expectations.
@@
Human Only Humans + PASTA
@@
PASTA’s increases expert agreement as its judgments are directionally aligned with human expert. When incorporated as a fourth evaluator in the intraclass correlation coefficients(ICC) calculations, PASTA further increased overall panel consistency, with ICC rising from .5940 to .6249 for violation score and from .6393 to .6607 for relevance score (see Table 6). According to ICC interpretation guidelines, these values represent moderate-to-good reliability, indicating that PASTA behaves more like a stable additional rater than a source of noise  (Braun and Clarke, 2006; Koo and Li, 2016).
@@
PASTA demonstrated stable and directionally consistent behavior across both violation and relevance scoring tasks. Its outputs closely tracked expert judgments in common, low-severity cases, while deviations in higher-severity violation scores followed a systematic conservative pattern—PASTA tended to slightly under-estimate severity rather than drift unpredictably. Relevance scoring spanned the full range and aligned more closely with expert distributions, reflecting that relevance judgments are easier for the model than detecting high-severity misalignment. Together, these patterns indicate that PASTA applies coherent internal criteria instead of producing noisy or erratic predictions, and that its conservative stance in high-severity cases reflects caution rather than rejection. Full confusion matrices and scatter plots illustrating these trends are provided in Appendix  C.
@@ 6. System Scalability: Runtime and Cost Evaluation
Ensuring scalability requires assessing both PASTA’s cost of running multi-policy compliance checks and its cost of incorporating new regulatory documents. Using 3 diverse model cards among the 8 used in the preprocessing procedure, we found that evaluating 5 major policies completes in approximately 1.5–1.8 minutes at a cost of \$2.86–\$3.06, demonstrating low and stable marginal cost as policies accumulate. In Section  6.2, we further show that adapting new policies incurs a similarly lightweight cost of only 2.24 minutes and \$4.44 even for a densely structured regulation like the EU AI Act. Together, these results indicate that PASTA can scale efficiently along both dimensions without imposing substantial computational overhead. LLM models used in both Section  6.1 and Section  6.2 are Claude Sonnet 4 and are ran on 2025-12-03.
@@ 6.1. Scalability of Multi-Policy Evaluation
Evaluation time shows linear–to–sublinear scaling, completing within 1.49–1.82 minutes across runs. Because policies are evaluated in parallel, total runtime is bounded by the longest single-policy evaluation rather than by the number of policies. For example, in run 3, all 5 policies were evaluated concurrently in 40.75 seconds (0.68 minutes), followed by 68.14 seconds for summary generation.
@@ 6.2. Overhead of New Policy Adaptation
PASTA’s scalability also depends on the cost of integrating new regulatory documents into the system as the current policy landscape is fast-evolving. Adapting a new policy requires two automated steps described in Section 4.4: (1) policy structuring, where the legal text is normalized into PASTA’s standardized table foramt, and (2) automated relevancy evaluation, where an LLM scores each comparison pair on a 0–5 relevance scale and constructs an irrelevancy map. For the EU AI Act, policy structuring completed in 94.88 seconds and cost \$3.48, processing all API requests concurrently. Automated relevancy evaluation required only 39.47 seconds and cost \$0.96. These results show that even for one of the longest and most structurally complex AI regulations, both steps incur minimal computational cost.
@@
Processing Step Time (min) Input Tokens Output Tokens Total Cost
@@ 7.1.2. Study Design
    Post-Task Survey (10 minutes): Participants completed the System Usability Scale (SUS), the NASA Task Load Index(NASTA-TLX), and Likert-scale items assessing both the model card process and the PASTA interface  (Brooke, 1995; Hart and Staveland, 1988).
@@ 7.2. Findings
Overall, PASTA proved usable (SUS $M=73.5$, $SD=15.2$) and effective in supporting multi-policy AI compliance evaluation. Participants described the model card input as effortful yet manageable, often framing it as a useful reflection on their systems. The output reports were seen as actionable and interpretable, and participants completed report reviews quickly, averaging 6.5 minutes ($SD=5.72$) when surfacing policy-specific issues and linking them to system components, though they wanted stronger guidance on remediation. The tool enabled participants to identify compliance risks quickly and with confidence. Figure. 7 summarizes participants’ responses to the usability questions, highlighting strong agreement on the tool’s clarity, actionability, and overall usefulness. We present these findings across 3 themes: the effort and value of input, the clarity and actionability of output, and the tool’s impact on practitioners’ overall understanding of compliance.
@@
Completing the model card functioned less like filling out a bureaucratic form and more like engaging in a structured reflection on the system. Although NASA-TLX ratings indicated consistently high mental demand (70–95; $M=76.7,SD=25.4$) and notable effort, participants framed this cognitive load as purposeful rather than burdensome. Instead of slowing them down, the card helped consolidate information they typically held in scattered notes, old IRB documents, or ad-hoc descriptions. As one participant put it, “It felt like a check-up on my system rather than random paperwork” (P3). Participants consistently described the model card as a productive prompt for articulating decisions, assumptions, and risks that had never been written down in one place. This suggests that the model card acted as a reflective scaffold.
The task was also relatively time-efficient compared to traditional compliance workflows and existing AI compliance tools, illustrating that the model card served as a structured shortcut rather than an added burden. Most participants completed the card in 20–30 minutes ($M=28.4$, $SD=6.7$; range 19–39.5), and several contrasted this with the multi-hour process of assembling documentation for legal consultations or governance reviews. For practitioners without access to enterprise governance tools or formal legal support, this consolidation of work was especially valuable. This comparatively low-effort input also stands out when benchmarked against existing compliance-oriented tools. Platforms offering “regulatory compliance checking” (such as LumeNova, VerifyWise, Collibra AI, and Credo AI; Section 2.1) rely on questionnaire-style inputs that request information similar in scope—system overviews, oversight mechanisms, misuse scenarios—but require this work to be repeated for each regulatory regime, causing effort to multiply as policies increase. Other tool categories offer no meaningful input-time baseline: enterprise governance platforms require multi-week technical integration; risk-prediction and policy-interpretation tools accept short free-text descriptions and do not perform compliance checks; and behavior-verification tools require source code or dataset access. Even model-card variants aligned with the EU AI Act, runtime governance, or human-centered communication do not report input-time measurements and target narrower goals. Notably, Tauqeer et al.’s GDPR knowledge-graph checker reports 3.55–11.48 minutes merely to encode a single contract instance for one policy. In contrast, PASTA’s $\sim$30-minute model-card input produces comprehensive, system-level documentation that supports multi-policy assessment (EU AI Act, AIDA, CCPA) in a single pass. Together, these comparisons indicate that PASTA offers a favorable effort-to-coverage ratio: it concentrates focused cognitive work into a reusable documentation artifact while avoiding the compounded burdens seen in existing solutions.
@@ 7.2.2. PASTA reorganized large multi-policy evaluations into decision-ready, component-level guidance
On the output side, PASTA largely achieved its goal of making large-scale, cross-jurisdiction evaluations interpretable and actionable (Design Goals 1 and 2). Participants reported strong confidence in their ability to surface problems across different policies ($M=4.73,SD=0.45$) and to see where risks were concentrated. Rather than treating policies as separate checklists, they used PASTA to build a comparative picture of deployment options. One participant explained, “I could check what problems might occur if I released a product in the U.S. That was very clear” (P2), while another described using the reports to reason about launch order: “It made it obvious that the EU is stricter; if resources are tight, we might start in California or Canada first” (P10). The tool thus helped translate a heterogeneous policy landscape into a configuration of practical choices.
PASTA also helped participants connect abstract policy clauses to specific parts of their systems, though support for remediation remained only partial. Survey results indicated high confidence in locating compliance issues within system components ($M=4.27$) and understanding why those issues arose ($M=4.18,SD=0.57$). Participants emphasized that the clause-linked issue descriptions “finally told me which piece of the pipeline each requirement was about” (P4). However, ratings for finding clear ways to fix issues were lower, though still moderate ($M=3.91,SD=1.16$). The issues and fixes table was frequently described as the most actionable element – “The issues and fixes table gave me concrete next steps…that’s what I really want to know: what do I need to fix?” (P7) – yet several participants asked for more structure in the recommendations, such as tagging fixes by effort level or affected subsystem. This pattern suggests that PASTA effectively supports sensemaking and diagnosis, while remediation planning remains an area for refinement.
@@
7.2.3. PASTA shifted practitioners’ conception of compliance from vague ethics to concrete, multi-policy obligations

### v2

@@ 4.6. Cost Optimization Strategies: Policy Chunking and Irrelevancy Filtering
Table 4. Empirical results showing how irrelevancy filtering reduces the number of comparison pairs and total runtime. The full evaluation compares all 23 model card sections against five policies–AIDA (40 articles), CCPA (54 articles), EU AI Act (113 articles), GDPR (99 articles), and the Colorado AI Act (8 articles)—resulting in 4,907 initial section–article pairs. Irrelevancy filtering removes pairs that consistently provide no contribution to compliance assessment, reducing the evaluation volume by 44.9% without loss of coverage.
@@
To support Design Goal 3, PASTA aggregates detailed pairwise evaluations into summaries that make large-scale compliance results easy to interpret. The evaluation process described in Section  4.5 generates a structured dataset consisting of one evaluation result for each comparison pair. Each result includes the corresponding model card section, policy article number, violation score, and an accompanying rationale. An example of the output is provided in Appendix.  10. The results are formatted as Markdown tables and are subsequently aggregated by the LLM into interpretable compliance summaries and actionable guidance. By condensing the detailed results into targeted summaries, the system helps users efficiently understand overall compliance status, identify policy-specific gaps, and highlight system components requiring attention. The evaluation results are aggregated into the following summaries:
  • Overall Summary: Provides a holistic assessment of the AI system’s compliance across all selected policies (see interface in Fig. 10 in Appendix  B.4).
  • Policy-wise Summary: Presents a focused evaluation of the system’s compliance with individual policies, offering policy-specific insights (see interface in Fig. 11 in Appendix  B.4).
@@
Figure 5. Section-wise Evaluation Interface linking compliance issues to specific model card sections. Each row presents the user’s documentation, detected issues, and concrete recommendations for remediation, supporting actionable system improvements. This table also supports policy-specific filtering.
@@ 5.1.1. Expert Recruitment
We recruited 3 AI–policy compliance experts from Upwork, a globally used freelancing platform commonly relied upon in academic research to source verified professional contractors with specialized expertise. Experts were selected based on demonstrated experience interpreting “AI regulatory frameworks,” and they were paid \$50 USD per hour. The 3 recruited experts reflected wide diversity in legal training. They were licensed attorneys, holding JD, LLB, and LLM degrees alongside jurisdiction-specific bar qualifications, with 6–10 years of legal practice across jurisdictions including the EU, North America, Latin America, the UK, and the Middle East. Their professional backgrounds covered privacy and data protection, AI governance and regulation, technology and cyberlaw, and commercial or corporate compliance. Each expert had prior experience assessing AI-related documentation and was familiar with major regimes such as the EU AI Act and CCPA.
@@ 5.1.2. Sampling of Policies and Model Cards
We sampled 18 policy articles across the EU AI Act and CCPA to ensure thematic breadth and balanced coverage, and sampled 3 diverse model cards among the 8 used in the preprocessing procedure detailed in Section  4.4 . To identify high-salience model card sections, we ran PASTA across all model-card sections and selected the 6 with the highest average violation scores across both policies. This resulted in 18 model-card sections which, when paired with the 18 curated articles, yielded 324 section–policy comparison pairs as our evaluation set. A full list of selected policy articles is provided in Appendix  B.
@@ 5.2. Findings
Overall, the results indicate that PASTA offers judgments that are comparably stable and human-aligned.
Before evaluating PASTA, we first assessed whether the 3 human experts formed a reliable evaluation baseline. Their intraclass correlation coefficients were .5940 for evaluating the violation score and .6393 for the relevance score, which can be interpreted as moderate-to-strong agreement based on established ICC guidelines  (Braun and Clarke, 2006; Koo and Li, 2016). This confirms that the expert-annotated dataset provides a consistent foundation for model comparison.
@@
PASTA’s compliance reasoning closely aligns with those of the human experts. Its violation and relevance scores closely mirrored expert results, reflected in Spearman correlations of $\rho=.626$ for violation scores and $\rho=.7611$ for relevance scores (see Table 5). Using widely cited interpretation thresholds (Cohen: .50–.69 = large; Mukaka: .60–.79 = strong), these values represent strong correlations  (Cohen, 1988; Mukaka, 2012). Its score magnitudes remained well-calibrated: mean absolute error was low (.284.284 for violation; .542.542 for relevance score), and the system matched expert scores within one point in 94.14% and 87.35% of cases respectively. These patterns indicate that PASTA approximates score levels in ways that align with expert expectations.
@@
Human Only Human + PASTA
@@
PASTA increases expert agreement as its judgments are directionally aligned with human judgements. When incorporated as a fourth evaluator in the intraclass correlation coefficients(ICC) calculations, PASTA further increased overall panel consistency, with ICC rising from .5940 to .6249 for violation score and from .6393 to .6607 for relevance score (see Table 6). According to ICC interpretation guidelines, these values represent moderate-to-good reliability, indicating that PASTA behaves more like a stable additional rater than a source of noise  (Braun and Clarke, 2006; Koo and Li, 2016).
@@
PASTA demonstrated stable and directionally consistent behavior across both violation and relevance scoring tasks. Its outputs closely tracked expert judgments in common, low-severity cases, while deviations in higher-severity violation scores followed a systematic conservative pattern—PASTA tended to slightly under-estimate severity rather than drift unpredictably. Relevance scoring spanned the full range and aligned more closely with expert distributions, reflecting that relevance judgments are easier for the model than detecting high-severity misalignment. Together, these patterns indicate that PASTA applies coherent internal criteria instead of producing noisy or erratic predictions, and that its conservative stance in high-severity cases reflects caution rather than rejection. Full confusion matrices and scatter plots illustrating these trends are provided in Appendix  B.3.
@@ 6. System Scalability: Runtime and Cost Evaluation
Ensuring scalability requires assessing both PASTA’s cost of running multi-policy compliance checks and its cost of incorporating new regulatory documents. Using 3 diverse model cards among the 8 used in the preprocessing procedure, we found that evaluating 5 major policies completes in approximately 1.5–1.8 minutes at a cost of \$2.86–\$3.06, demonstrating low and stable marginal cost as policies accumulate. In Section  6.2, we further show that adapting new policies incurs a similarly lightweight cost of only 2.24 minutes and \$4.44 even for a densely structured regulation like the EU AI Act. Together, these results indicate that PASTA can scale efficiently along both dimensions without imposing substantial computational overhead. LLM models used in both Section  6.1 and Section  6.2 are Claude Sonnet 4 and were ran on 2025-12-03.
@@ 6.1. Scalability of Multi-Policy Evaluation
Evaluation time shows linear to sublinear scaling, completing within 1.49–1.82 minutes across runs. Because policies are evaluated in parallel, total runtime is bounded by the longest single-policy evaluation rather than by the number of policies. For example, in run 3, all 5 policies were evaluated concurrently in 40.75 seconds (0.68 minutes), followed by 68.14 seconds for summary generation.
@@ 6.2. Overhead of New Policy Adaptation
PASTA’s scalability also depends on the cost of integrating new regulatory documents into the system as the current policy landscape is fast-evolving. Adapting a new policy requires two automated steps described in Section 4.4: (1) policy structuring, where the legal text is normalized into PASTA’s standardized table format, and (2) automated relevancy evaluation, where an LLM scores each comparison pair on a 0–5 relevance scale and constructs an irrelevancy map. For the EU AI Act, policy structuring completed in 94.88 seconds and cost \$3.48, processing all API requests concurrently. Automated relevancy evaluation required only 39.47 seconds and cost \$0.96. These results show that even for one of the longest and most structurally complex AI regulations, both steps incur minimal computational cost.
@@
Processing Step Time Input Output Cost
(min) Tokens Tokens
@@ 7.1.2. Study Design
    Post-Task Survey (10 minutes): Participants completed the System Usability Scale (SUS), the NASA Task Load Index (NASA-TLX), and Likert-scale items assessing both the model card process and the PASTA interface  (Brooke, 1995; Hart and Staveland, 1988).
@@ 7.2. Findings
Overall, PASTA proved usable (SUS $M=73.5$, $SD=15.2$) and effective in supporting multi-policy AI compliance evaluation. Participants described the model card input as effortful yet manageable, often framing it as a useful reflection on their systems. The output reports were seen as actionable and interpretable, and participants completed report reviews quickly, averaging 6.5 minutes ($SD=5.72$) when surfacing policy-specific issues and linking them to system components, though they wanted stronger guidance on remediation. The tool enabled participants to identify compliance risks quickly and with confidence. Fig.  7 summarizes participants’ responses to the usability questions, highlighting strong agreement on the tool’s clarity, actionability, and overall usefulness. We present these findings across three themes: the effort and value of input, the clarity and actionability of output, and the tool’s impact on practitioners’ overall understanding of compliance.
@@
Completing the model card functioned less like filling out a bureaucratic form and more like engaging in a structured reflection on the system. Although NASA-TLX ratings indicated consistently high mental demand (70–95; $M=76.7,SD=25.4$) and notable effort, participants framed this cognitive load as purposeful rather than burdensome. Instead of slowing them down, the card helped consolidate information they typically held in scattered notes, old IRB documents, or ad–hoc descriptions. As one participant put it, “It felt like a check-up on my system rather than random paperwork” (P3). Participants consistently described the model card as a productive prompt for articulating decisions, assumptions, and risks that had never been written down in one place. This suggests that the model card acted as a reflective scaffold.
The task was also relatively time-efficient compared to traditional compliance workflows and existing AI compliance tools, illustrating that the model card served as a structured shortcut rather than an added burden. Most participants completed the card in 20–30 minutes ($M=28.4$, $SD=6.7$; range 19–39.5), and several contrasted this with the multi-hour process of assembling documentation for legal consultations or governance reviews. For practitioners without access to enterprise governance tools or formal legal support, this consolidation of work was especially valuable. This comparatively low-effort input also stands out when benchmarked against existing compliance-oriented tools. Platforms offering “regulatory compliance checking” (such as LumeNova, VerifyWise, Collibra AI, and Credo AI; Section  2.1) rely on questionnaire–style inputs that request information similar in scope—system overviews, oversight mechanisms, misuse scenarios—but require this work to be repeated for each regulatory regime, causing effort to multiply as policies increase. Other tool categories offer no meaningful input-time baseline: enterprise governance platforms require multi-week technical integration; risk-prediction and policy-interpretation tools accept short free-text descriptions and do not perform compliance checks; and behavior-verification tools require source code or dataset access. Even model-card variants aligned with the EU AI Act, runtime governance, or human-centered communication do not report input–time measurements and target narrower goals. Notably, Tauqeer et al.’s GDPR knowledge-graph checker reports 3.55–11.48 minutes merely to encode a single contract instance for one policy. In contrast, PASTA’s $\sim$30-minute model-card input produces comprehensive, system-level documentation that supports multi-policy assessment (EU AI Act, AIDA, CCPA) in a single pass. Together, these comparisons indicate that PASTA offers a favorable effort-to-coverage ratio: it concentrates focused cognitive work into a reusable documentation artifact while avoiding the compounded burdens seen in existing solutions.
@@ 7.2.2. PASTA reorganized large multi-policy evaluations into decision-ready, component-level guidance
On the output side, PASTA largely achieved its goal of making large–scale, cross–jurisdiction evaluations interpretable and actionable (Design Goals 1 and 2). Participants reported strong confidence in their ability to surface problems across different policies ($M=4.73,SD=0.45$) and to see where risks were concentrated. Rather than treating policies as separate checklists, they used PASTA to build a comparative picture of deployment options. One participant explained, “I could check what problems might occur if I released a product in the U.S. That was very clear” (P2), while another described using the reports to reason about launch order: “It made it obvious that the EU is stricter; if resources are tight, we might start in California or Canada first” (P10). The tool thus helped translate a heterogeneous policy landscape into a configuration of practical choices.
PASTA also helped participants connect abstract policy clauses to specific parts of their systems, though support for remediation remained only partial. Survey results indicated high confidence in locating compliance issues within system components ($M=4.27$) and understanding why those issues arose ($M=4.18,SD=0.57$). Participants emphasized that the clause-linked issue descriptions “finally told me which piece of the pipeline each requirement was about” (P4). However, ratings for finding clear ways to fix issues were lower, though still moderate ($M=3.91,SD=1.16$). The issues-and-fixes table was frequently described as the most actionable element – “The issues-and-fixes table gave me concrete next steps…that’s what I really want to know: what do I need to fix?” (P7) – yet several participants asked for more structure in the recommendations, such as tagging fixes by effort level or affected subsystem. This pattern suggests that PASTA effectively supports sensemaking and diagnosis, while remediation planning remains an area for refinement.
@@
7.2.3. PASTA shifted practitioners’ conception of compliance from vague ethics to concrete, multi–policy obligations